A woman on a bridge 74.1 m high sees a raft floating at a constant speed on the river below. She drops a stone from rest in an attempt to hit the raft. The stone is released when the raft has 7.13 m more to travel before passing under the bridge. The stone hits the water 3.91 m in front of the raft. Calculate the speed of the raft.
Answer #1

motion of stone ::

initial speed = Vo = 0 m/s

acceleration = a = 9.8 m/s2

t = time taken

d = displacement = height = 74.1 m

Using the equation

d = Vo t + (0.5) a t2

74.1 = 0 x t + (0.5) (9.8) t2

t = 3.89 sec

Motion of raft :

Speed = V

distance travelled = 7.13 - 3.91 = 3.22 m

time = t = 3.89 sec

Using the formula

distance = V t

V = 3.22 / 3.89 = 0.83 m/s